2. Find the solution of the second order differential equations: day + y = 0, y(TT/3) = 0, y(TT/3) = dx2 a. = 4 b. y – 8y
Answer #1

Now the complementary function is given by Solution: y (11/3) = 4 2 (a) d²y y (7/3)=0 ty daz Now axillayny equation, mat i =Get thene G = -√3 (2 =-213 Hence the solution is. YCR)= -2/5 cosn +2 sina y (2) = a sina - 285 cosa # (6 Solution :- Ciren, yNow the final general solution to (0) = c e° о.с. = 1 Ас = 1 76) — 5 90) = 4 се + с. (че“e) = о - 4e14 ulse) = 0 ) Cse =-4